# 2HandedAfforder: Learning Precise Actionable Bimanual Affordances from Human Videos

### v2

@@ Abstract
When interacting with objects, humans effectively reason about which regions of objects are viable for an intended action, i.e., the affordance regions of the object. They can also account for subtle differences in object regions based on the task to be performed and whether one or two hands need to be used. However, current vision-based affordance prediction methods often reduce the problem to naive object part segmentation. In this work, we propose a framework for extracting affordance data from human activity video datasets. Our extracted 2HANDS dataset contains precise object affordance region segmentations and affordance class-labels as narrations of the activity performed. The data also accounts for bimanual actions, i.e., two hands co-ordinating and interacting with one or more objects. We present a VLM-based affordance prediction model, 2HandedAfforder, trained on the dataset and demonstrate superior performance over baselines in affordance region segmentation for various activities. Finally, we show that our predicted affordance regions are actionable, i.e., can be used by an agent performing a task, through demonstration in robotic manipulation scenarios.
@@ 1 Introduction
Figure 1: A motivating example: When required to label affordances for a task ‘Pour into bowl’, the typically labeled affordances provided by annotators are not precise and reduce the problem to object part segmentation. Alternatively, our affordance extraction method uses the hand-object interaction sequence to obtain precise, actionable, bimanual object affordance regions.
@@
We argue that affordances should not be labeled but automatically extracted from observations of humans performing tasks, such as human activity video datasets. We propose a method that uses hand-inpainting and mask completion to extract affordance regions occluded by human hands. This has several advantages. First, by using this procedure, we are able to obtain bimanual and precise affordances (Figure  1) rather than simply predicting object parts. Second, it makes affordance specification more natural since it is often easier for humans to show the object region to interact with, rather than label and segment it correctly in an image. Third, using human activity videos gives us diverse task-specific affordances, with the affordance class label naturally coming from the narration of what task is being done by the human. This makes our affordances task-oriented with natural language specification, unlike previous methods focused on predicting task-agnostic interaction hotspots [10, 1].
We extract a dataset, 2HANDS (2-Handed Affordance + Narration DataSet), consisting of a large number of unimanual and bimanual object affordance segmentation masks and task narrations as affordance class-labels. We propose a VLM-based affordance prediction model, 2HandedAfforder, that is trained on the 2HANDS dataset and predicts affordance masks in images based on an input text prompt. To evaluate the performance on this challenging problem, we also present a novel benchmark, ActAffordance, using annotations on images from two egocentric human activity datasets [5, 11].
Our contributions are:
@@
  • a novel benchmark, ActAffordance, to evaluate affordance detection aligned with the ”ground truth” considered by humans.
@@ 2 Related Work
Our approach and goals. In this work, we propose a method to extract affordance masks leveraging recent video-based hand inpainting techniques [2]. Since our dataset contains precise segmentation masks, we can predict pixel-wise affordance segments in the image, as opposed to methods only trained with point-labels of affordance [35] or that only predict heatmaps [28, 1]. Moreover, we consider the especially challenging problem of bimanual affordance detection, for which the spatial context of the objects and their interconnection is also important. Although bimanual affordances have been considered in previous work [19, 9, 42, 33], to the best of our knowledge, ours is the first method to extract bimanual affordances from videos which we then use to train our model to predict task-specific affordance masks based on a text prompt.
@@ 3.1 Affordance Extraction from Human Videos
With the hand and object masks available over the entire video sequence, we obtain an un-occluded view of the objects by inpainting out the hands. We use a video-based hand inpainting model, VIDM [2], that uses 4 frames from the sequence as input to inpaint the missing regions. This makes reconstructing the target objects more likely if the objects have already been in another frame of the action sequence without occlusion. Inpainting provides us with an un-occluded view of the objects. We then precisely segment these un-occluded objects in the inpainted image using mask completion. For this, we use the segmentation masks from the original image and prompt SAM2 [39] to propagate these masks to the new inpainted image. We observe that this process gives us more precise object masks compared to directly using mask completion methods [43], detailed in the appendix (Sec. 12).
@@ 4.1 ActAffordance Benchmark
To answer the first question of the accuracy of our extracted affordances in the 2HANDS dataset, we evaluate the alignment of our extracted affordance masks with human-annotated affordance regions. As mentioned in Sec. 3.1, when humans label affordances, they often simply label object parts and do not necessarily focus on the precise regions of interaction of the objects [28, 35]. Moreover, the second question regarding the accuracy of 2HandedAfforder is non-trivial. Using only the masks in our 2HANDS dataset as “ground truth” leads to a bias towards our extracted affordances. This may not be in alignment with what humans consider accurate affordance regions. Therefore, we propose a novel benchmark called “ActAffordance” to evaluate both the dataset quality and the predicted affordances.
@@
We note that ActAffordance is a particularly challenging benchmark. To date, reasoning segmentation, i.e., text- prompt-based segmentation of full objects, is an unsolved problem. Prompt-based segmentation of precise object affordance regions is yet more challenging, especially when benchmarked against humans. The addition of bimanual affordances with multiple objects is another step beyond that. However, we feel this challenging benchmark will push the community forward towards building more effective methods for affordance prediction and thus we evaluate all methods and baselines on this benchmark instead of a simpler test set from our dataset.
4.2 Metrics for Evaluation
We use several metrics to evaluate the performance of the proposed models and baselines. Since we treat the affordance detection problem as a segmentation task, we use metrics widely spread across segmentation problems: precision, intersection over union (IoU), and the directed and general Hausdorff Distance (HD). We train our 2HandedAfforder and 2HandedAfforder-CLIP models on the 2HANDS dataset and evaluate on the ”ActAffordance” benchmark. We evaluate performance on both the EPIC-KITCHENS and Ego4D splits of the benchmark. Note that there is no overlap between the data from the EPIC-KITCHENS data used in 2HANDS. The evaluation on the Ego4D split of the benchmark also helps us answer the generalization question since there is no Ego4D data in 2HANDS.
Note that for the evaluation of our models, the false negative predictions play a reduced role since our models are not trained to predict all of the multimodal solutions in the benchmark but to predict precise affordance regions which might only cover a subset of all of the possible solutions. Thus, the key metric for comparison is precision over IoU. Another common metric for segmentation is the Hausdorff distance (HD). For each point in each set, the distance to the closest point from the other set is calculated and the Hausdorff distance is defined as the maximum of all of these distances. Similar to the IoU, including the distance from the ground truth to the prediction might distort the results since we aim to predict precise affordances that may only cover a smaller subset of the ground truth. Thus, we additionally provide the directed Hausdorff distance that only calculates the maximum distance from the prediction set points to the ground truth set.
@@ 4.2 Metrics for Evaluation
Figure 5: Qualitative affordance prediction results on the ActAffordance benchmark. We compare our 2HandedAfforder model against LISA [20], AffordanceLLM [37] and 3DOI [35]. We also include an example result if we were to run our affordance extraction method on the activity sequence to show the quality of the extraction. Red and green masks denote the left hand and right hand affordance mask predictions, respectively.
@@
5 Results
5.1 Affordance Extraction Quality
We assess the quality of the affordances obtained from our extraction pipeline (Sec. 3.1) by evaluating their alignment with the human annotations in the ActAffordance benchmark. The comparison results are shown in Table 2, “AffExtract”, and Figure 4.2. As noted previously, the benchmark annotations contain all the possible modes of object interaction, while the extraction process and our models only cover a single interaction mode. Thus, precision is the most important metric to evaluate over IoU. The same principle is true for the Hausdorff distance (HD), which is why we also report Directional Hausdorff distance (Dir. HD), which only calculates the maximum distance from the prediction set points to the ground truth set. We note the precision of AffExtract is better for the Ego4D split (0.541) than the EPIC-KITCHENS split (0.334) with a combined score of 0.42. This shows a reasonably good alignment with the human-annotated segmentations from the benchmark and meaningful affordance region extraction. The IoU scores are relatively lower, with an average of 0.185, showing the challenge of the task when compared against human-level object understanding.
@@ 5.1 Affordance Extraction Quality
Figure 6: Examples of different manipulation tasks executed on a bimanual Tiago++ robot. Red and green masks denote left and right hand affordance mask predictions, respectively. We segment the task-specific object affordance regions, propose grasps for these regions, and use pre-designed motion primitives to execute manipulation tasks. Videos at: https://sites.google.com/view/2handedafforder.
@@ 5.2 Comparison against baselines on ActAffordance benchmark
All models were evaluated on the ActAffordance benchmark. Additionally, we assessed the methods on a modified version of the benchmark, where all images were cropped to encompass the target objects to enhance comparability with baselines that utilize bounding boxes [1] or query points [35] as prompts instead of language. Since LOCATE [22] uses an explicit affordance class label as input for prediction, we adapt the EPIC-VISOR verb categories used in 2HANDS to fit the AGD20K classes used in LOCATE. Such an adaptation is not possible for Ego4D so we exclude LOCATE from the comparison on the Ego4D split.
@@ 5.3 In-the-wild Affordance Prediction and Robot Demonstration
We conduct robotic manipulation experiments with various objects using a bimanual Tiago++ robot in a realistic kitchen environment.We deployed our best-performing 2HandedAfforder model for affordance region segmentation inference based on user-defined task prompts such as ‘pour into cup’.
@@
We demonstrate how our affordance prediction method improves the performance of a robot in executing manipulation tasks compared to using standard object or part segmentation approaches, such as the mask output of LangSAM. By integrating our affordance prediction into the grasping pipeline, the robot is able to make more informed grasping decisions, leading to greater task success. Examples of different manipulation tasks are shown in Figure 6, and videos of the execution are provided at https://sites.google.com/view/2handedafforder.
@@ 6 Conclusion
In this work, we proposed a framework for extracting precise, meaningful affordance regions from human activity videos, resulting in the 2HANDS dataset of actionable bimanual affordances. We further introduced a novel VLM-based task-aware bimanual affordance prediction model, 2HandedAfforder, that predicts actionable affordance regions from task-related text prompts. To evaluate the alignment of the extracted affordances with human-annotated ones, we further proposed a novel ActAffordance benchmark, which is a particularly challenging benchmark for prompt-based segmentation of precise object affordance regions. Our experiments demonstrate that 2HandedAfforder can predict meaningful task-oriented bimanual affordances compared to other works, thereby showcasing the effectiveness of our data extraction pipeline and proposed model. Our real-world robotic manipulation experiments further showcased the efficacy of the predicted affordances compared to other object or part segmentation methods. In the future, we plan to extend our method to other egocentric video datasets and to other tasks beyond the kitchen context. We also plan to use activity recognition methods to automatically extract narrations for the videos.
@@ 12 Qualitative Analysis of Mask Completion Approaches
We developed and evaluated two mask completion approaches, i.e., an image reconstruction (IR) based and video segmentation (VS) based approach. The IR based approach uses the image inpainting model MI-GAN [43] to inpaint the missing regions of the mask using the hand mask as inpainting region. The VS based approach creates an image sequence out of the original image and the image with the hands removed. The object mask for the original image is then propagated to the inpainted image using SAM2 [39] to create the completed version of the mask. The evaluation of these approaches was conducted qualitatively, and some examples can be seen in Fig. 9. It is clearly visible that the VS based approach performs better on average than the IR based approach. Generally, the VS based approach provides more accurate results, see row 1 and 2, and it does not detect affordances at regions where the object was not reconstructed properly (row 3 and 4). This can be explained intuitively for two reasons: The IR based approach has no information about the underlying RGB image and only processes the object mask itself which is binary by nature. So the image reconstruction model focuses on simple principles such as the continuation of lines and shapes. This leads to the reconstruction of what we call ’ghost handles’ where the image reconstruction model still predicts the existence of a handle or object part in general even though the object was not reconstructed successfully. This also reduces the accuracy of the IR based method. The VS based approach, however, has the information about the inpainted image and thus will only complete object parts that are properly inpainted. So, it naturally filters out data points where the object was not inpainted properly since the completed mask will not intersect the hand mask. Thus, it will always be more accurate and never predict ’ghost handles’. There are only a couple of examples where the IR based approach performs better than the VS based approach (row 5). Hence, we decided to use the VS based approach for the creation of 2HANDS.

### v3

@@ Abstract
When interacting with objects, humans effectively reason about which regions of objects are viable for an intended action, i.e., the affordance regions of the object. They can also account for subtle differences in object regions based on the task to be performed and whether one or two hands need to be used. However, current vision-based affordance prediction methods often reduce the problem to naive object part segmentation. In this work, we propose a framework for extracting affordance data from human activity video datasets. Our extracted 2HANDS dataset contains precise object affordance region segmentations and affordance class-labels as narrations of the activity performed. The data also accounts for bimanual actions, i.e., two hands co-ordinating and interacting with one or more objects. We present a VLM-based affordance prediction model, 2HandedAfforder, trained on the dataset and demonstrate superior performance over baselines in affordance region segmentation for various activities. Finally, we show that our predicted affordance regions are actionable, i.e., can be used by an agent performing a task, through demonstration in robotic manipulation scenarios. Project-website: sites.google.com/view/2handedafforder
@@ 1 Introduction
Figure 1: A motivating example: When labeling affordances for a task ‘Pour into bowl’, typical labeled affordances provided by annotators are not precise and reduce the problem to object part segmentation. Alternatively, our affordance extraction method uses the hand-object interaction sequence to get precise bimanual affordance regions that are not just ‘conceptual’ but also ‘actionable’.
@@
We argue that affordances should not be labeled but automatically extracted by observing humans performing tasks, e.g. in activity video datasets. We propose a method that uses hand-inpainting and mask completion to extract affordance regions occluded by human hands. This has several advantages. First, by using this procedure, we are able to obtain bimanual and precise affordances (Figure  1) rather than simply predicting object parts. Second, it makes affordance specification more natural since it is often easier for humans to show the object region to interact with, rather than label and segment it correctly in an image. Third, using human activity videos gives us diverse task-specific affordances, with the affordance class label naturally coming from the narration of what task is being done by the human. This makes our affordances task-oriented with natural language specification, unlike previous methods focused on predicting task-agnostic interaction hotspots [12, 1].
We extract a dataset, 2HANDS (2-Handed Affordance + Narration DataSet), consisting of a large number of unimanual and bimanual object affordance segmentation masks and task narrations as affordance class-labels. We propose a VLM-based affordance prediction model, 2HandedAfforder, that is trained on the 2HANDS dataset and predicts affordance masks in images based on an input text prompt. To evaluate the performance on this challenging problem, we also present a novel benchmark, ActAffordance, using annotations on images from two egocentric human activity datasets [5, 13]. Our contributions are:
@@
  • a new benchmark, ActAffordance, with affordance annotations by humans who observe the interaction sequence.
@@ 2 Related Work
Our approach and goals. In this work, we propose a method to extract affordance masks leveraging recent video-based hand inpainting techniques [2]. Since our dataset contains precise segmentation masks, we can predict pixel-wise affordance segments in the image, as opposed to methods only trained with point-labels of affordance [38] or that only predict heatmaps [31, 1, 8]. Moreover, we consider the especially challenging problem of bimanual affordance detection, for which the spatial context of the objects and their interconnection is also important. Although bimanual affordances have been considered in previous work [21, 11, 45, 30, 36, 9], to the best of our knowledge, ours is the first method to extract bimanual affordances from videos which we then use to train our model to predict task-specific affordance masks based on a text prompt.
@@ 3 Extraction and Learning of Bimanual Affordances from Human Videos
In this section, we detail our affordance extraction approach used to generate our 2HANDS dataset from videos of humans performing everyday tasks (Sec. 3.1). Then, we present our approach, “2HandedAfforder”, for predicting meaningful task-oriented bimanual affordance regions in images in Sec. 3.2.
@@ 3.1 Affordance Extraction from Human Videos
With the hand and object masks available over the entire video sequence, we obtain an un-occluded view of the objects by inpainting out the hands. We use a video-based hand inpainting model, VIDM [2], that uses 4 frames from the sequence as input to inpaint the missing regions. This sequence-based inpainting better reconstructs the target objects since the objects may be visible in another frame of the sequence without occlusion. Inpainting provides us with an un-occluded view of the objects. We then precisely segment these un-occluded objects in the inpainted image using mask completion. For this, we use the segmentation masks from the original image and prompt SAM2 [42] to propagate these masks to the new inpainted image. We observe that this process gives us more precise object masks compared to directly using mask completion methods [46], detailed in the appendix (Sec. 12).
@@ 4.1 ActAffordance Benchmark
To answer the first question of the accuracy of our extracted affordances in the 2HANDS dataset, we evaluate the alignment of our extracted affordance masks with human-annotated affordance regions. As mentioned in Sec. 3.1, when humans label affordances, they often simply label object parts and do not necessarily focus on the precise regions of interaction of the objects [31, 38]. Moreover, the second question regarding the accuracy of 2HandedAfforder is non-trivial. Using only the masks in our 2HANDS dataset as “ground truth” leads to a bias towards our own extracted affordances. Therefore, we propose a novel benchmark called “ActAffordance” to evaluate both the dataset quality and the predicted affordances. Specifically, we evaluate the alignment of our affordances with the affordances annotated by humans who are shown the interaction video sequence.
@@
We note that ActAffordance is a very challenging benchmark. To date, reasoning segmentation, i.e., text-prompt-based segmentation of full objects, is an unsolved problem. Prompt-based segmentation of precise object affordance regions is yet more challenging, especially when benchmarked against humans. The inclusion of bimanual affordances with multiple objects is another step beyond that. However, we feel this challenging benchmark will push the community forward towards more effective affordance prediction and thus we evaluate all methods on this benchmark instead of directly using the test set from our dataset.
@@
LOCATE Refer to caption Refer to caption Refer to caption
VRB Refer to caption Refer to caption Refer to caption
@@
Figure 5: Qualitative affordance prediction results on the ActAffordance benchmark. We compare our 2HandedAfforder model against LOCATE [24], VRB [1], LISA [22], AffordanceLLM [40], 3DOI [38]. We also include an example result if we run our affordance extraction method on the activity sequence to show the quality of the extraction. Red and green masks denote left and right hand affordance mask predictions, respectively.
@@
4.2 Metrics for Evaluation
Since we treat the affordance detection problem as a segmentation task, we use the following metrics to evaluate the performance of the proposed models and baselines: precision, intersection over union (IoU) and the directed and general Hausdorff distance (HD). We train our 2HandedAfforder and 2HandedAfforder-CLIP models on the 2HANDS dataset and evaluate on the “ActAffordance” benchmark. We evaluate performance on both the EPIC-KITCHENS and Ego4D splits of the benchmark. Note that there is no overlap between the data from EPIC-KITCHENS used in 2HANDS. The evaluation on the Ego4D split of the benchmark also helps answer the generalization question since there is no Ego4D data in 2HANDS.
Note that for the evaluation of our models, false negative predictions play a reduced role since our models are not trained to predict all multimodal solutions in the benchmark but to predict precise affordance regions which might only cover a subset of the possible solutions. Thus, the key metric for comparison is precision over IoU. Another common segmentation metric is Hausdorff distance (HD). For each point in each set, the distance to the closest point from the other set is computed and the Hausdorff distance is the maximum of all of these distances. Similar to the IoU case, including the distance from ground truth to the prediction might distort the results since we aim to predict precise affordances that may only cover a smaller subset of the ground truth. Thus, we also provide the directed Hausdorff distance that only calculates the maximum distance from the prediction set points to the ground truth set.
@@ 4.2 Metrics for Evaluation
Figure 6: Examples of different manipulation tasks executed on a bimanual Tiago++ robot. Red and green masks denote left and right hand affordance mask predictions, respectively. We segment the task-specific object affordance regions, propose grasps for these regions, and use pre-designed motion primitives to execute manipulation tasks. Videos are available at sites.google.com/view/2handedafforder.
5 Results
5.1 Affordance Extraction Quality
We assess the quality of affordances obtained from our extraction pipeline (Sec. 3.1) by evaluating their alignment with the human annotations in the ActAffordance benchmark. The results are shown in Table 2, “AffExtract”, and Figure 4.1. As noted before, the benchmark annotations contain all the possible modes of object interaction, while the extraction process and our models only cover a single interaction mode. Thus, precision is the most important metric to evaluate over IoU. The same principle is true for the Hausdorff distance (HD), which is why we also report directional Hausdorff distance (Dir. HD), which only calculates the maximum distance from the prediction set points to the ground truth set. We note the precision of AffExtract is better for the Ego4D split (0.541) than the EPIC-KITCHENS split (0.334) with a combined score of 0.42. This shows a reasonably good alignment with the human-annotated segmentations from the benchmark and meaningful affordance region extraction. The IoU scores are relatively lower, with an average of 0.185, showing the challenge of the task when compared against human-level object understanding.
@@ 5.2 Comparison against baselines on ActAffordance benchmark
All models are evaluated on the ActAffordance benchmark. Additionally, we assess the methods on a modified version of the benchmark, where all images were cropped to encompass the target objects for comparison with VRB that utilizes bounding boxes [1] and 3DOI that uses query points [38] as prompts instead of language. Although no baseline can be trained on our 2HANDS dataset, we make several adjustments to allow inference on the benchmark. For LISA, LOCATE, AffLLM, we ignore left/right classification and compare predicted masks to the union of left and right masks in the benchmark. For VRB, 3DOI, we input the necessary ground-truth bounding boxes and object mask centers (cropped benchmark) and predict separate left/right masks. Since LOCATE [24] uses an explicit affordance class label as input, we adapt the EPIC-VISOR verb categories used in 2HANDS to fit the AGD20K classes used in LOCATE. Such an adaptation is not possible for Ego4D so we exclude LOCATE from the comparison on the Ego4D split. To isolate the effect of the 2HANDS dataset, the comparison with AffLLM and LISA is key since their network architecture is close to ours.
@@ 5.3 In-the-wild Affordance Prediction and Robot Demonstration
We conduct robotic manipulation experiments with various objects using a bimanual Tiago++ robot in a realistic kitchen environment. We deploy our 2HandedAfforder model for affordance region segmentation inference based on task prompts such as ‘pour into cup’.
@@
We demonstrate how our affordance prediction method improves the performance of a robot in executing manipulation tasks compared to using standard object or part segmentation approaches, such as the mask output of LangSAM. By integrating our affordance prediction into the grasping pipeline, the robot is able to make more informed grasping decisions, leading to greater task success. Examples of different manipulation tasks are shown in Figure 6 and in videos at sites.google.com/view/2handedafforder.
@@ 6 Conclusion
In this work, we proposed a framework for extracting precise, meaningful affordance regions from human activity videos, resulting in the 2HANDS dataset of actionable bimanual affordances. We further introduced a novel VLM-based task-aware bimanual affordance prediction model, 2HandedAfforder, that predicts actionable affordance regions from task-related text prompts. To evaluate the alignment of the extracted affordances with human-annotated ones, we further proposed a novel ActAffordance benchmark, which is a particularly challenging benchmark for prompt-based segmentation of precise object affordance regions. Our experiments demonstrate that 2HandedAfforder can predict meaningful task-oriented bimanual affordances compared to other works, thereby showcasing the effectiveness of our data extraction pipeline and proposed model.
@@ 7 Additional Filtering and Augmentation Steps
The full dataset and codebase will be released at sites.google.com/view/2handedafforder.
@@ 12 Qualitative Analysis of Mask Completion Approaches
We developed and evaluated two mask completion approaches, i.e., an image reconstruction (IR) based and video segmentation (VS) based approach. The IR based approach uses the image inpainting model MI-GAN [46] to inpaint the missing regions of the mask using the hand mask as inpainting region. The VS based approach creates an image sequence out of the original image and the image with the hands removed. The object mask for the original image is then propagated to the inpainted image using SAM2 [42] to create the completed version of the mask. The evaluation of these approaches was conducted qualitatively, and some examples can be seen in Fig. 11. It is clearly visible that the VS based approach performs better on average than the IR based approach. Generally, the VS based approach provides more accurate results, see row 1 and 2, and it does not detect affordances at regions where the object was not reconstructed properly (row 3 and 4). This can be explained intuitively for two reasons: The IR based approach has no information about the underlying RGB image and only processes the object mask itself which is binary by nature. So the image reconstruction model focuses on simple principles such as the continuation of lines and shapes. This leads to the reconstruction of what we call ‘ghost handles’ where the image reconstruction model still predicts the existence of a handle or object part in general even though the object was not reconstructed successfully. This also reduces the accuracy of the IR based method. The VS based approach, however, has the information about the inpainted image and thus will only complete object parts that are properly inpainted. So, it naturally filters out data points where the object was not inpainted properly since the completed mask will not intersect the hand mask. Thus, it will always be more accurate and never predict ‘ghost handles’. There are only a couple of examples where the IR based approach performs better than the VS based approach (row 5). Hence, we decided to use the VS based approach for the creation of 2HANDS.